The Length of a rectangular flower bed is 3 times the width. The area of the bed is 108 square meters. What are the dimensions of the bed?
Please Help me. Thank You

Answer by checkley79(3341)

L=3W
AREA=LW
108=3W*W
108=3W^2
W^2=108/3
W^2=36
W=SQRT36
W=6 ANS. FOR THE WIDTH.
L=3*6=18 ANS. FOR THE LENGTH.
PROOF:
108=18*6
108=108